OpenStudy (annie12m):

Variable y varies directly with variable x, and y = 4 when x = 7. What is the value of x when y = 12?

OpenStudy (annie12m):

@silveralchemist09

OpenStudy (anonymous):

okay like before y=kx since it is directly varying now put values of y and x to find k which is constant of proportionality 4=k7 4/7=k now y=4/7x 12=4/7x x=21 ans

OpenStudy (anonymous):

Y=kx 4=7k k=4/7 12=4/7x 12*7/4=x x=21